OpenStudy (ttop0816):

Given the system of equations, what is the value of the system determinant? 2x + y = 8 x - y = 10 -1 -2 -3

OpenStudy (the_fizicx99):

The y's cancel out.. So its 2x + x = 8 + 10 3x = 18 /3 /3 x = 6

OpenStudy (the_fizicx99):

Oh, matrix? 2 + 1 1 - 1 = 2*(-1) - 1*1 = -2 - 1 = -3